OpenStudy (anonymous):

A moving company is trying to store boxes in a storage room with a length of 5 m, width of 3 m and height of 2 m. How many boxes can fit in this space if each is 10 cm long, 6 cm wide and 4 cm high?

OpenStudy (anonymous):

So you are given dimensions of a room, and dimensions of a box Volume of rectangle = lwh Divide volume of room by volume of box to find how much it can hold, be careful with units

OpenStudy (richyw):

ok first of all you can't always just divide the volume of the room by the volume of the boxes, because you will not always be able to fill the room completely with boxes. In this case however it is easy to recognize that the boxes will all fit completely since \[\frac{l_ {room}}{l_{box}},\frac{w_ {room}}{w_{box}},\frac{h_ {room}}{h_{box}}\]are all integer values. So now you could do the volume thing, but there is no need to because you will also notice that those rations described above are all 50. so your answer will simply be \(50^3\)

OpenStudy (anonymous):

yea but typically in math problems, it's not a real-life scenario, so you kind of have to assume the room is completely filled